# bibfile.bib

### v1

@@ 2.2 Existence of harmonic functions
An approach with harmonic functions was also used in [duc2015sup], where Duchesne proved a superrigidity result for cocompact lattices in higher rank semisimple Lie groups acting on spaces in the family $\mathcal{X}$.
Recall that $\Gamma$ is a discrete, torsion-free, cocompact subgroup of $G$. Its action on $\Sigma$ is free and proper ([garland], Lemma 2.6). Therefore, $\Gamma$ is the fundamental group of the finite simplicial complex $\sigma=\Sigma/\Gamma$, and $\Gamma$-equivariant functions $f:\Sigma(0)\rightarrow X$ correspond to functions $f_{\sigma}:\sigma(0)\rightarrow X$, where $\sigma(0)$ is the set of vertices in $\sigma$. The correspondence is defined by choosing a fundamental domain $D$ of $\Gamma$ in $\Sigma$ and assigning to $f_{\sigma}$ the values that $f$ takes at the vertices in $D$. Since $f$ is $\Gamma$-equivariant, all its values are uniquely determined by those at the vertices in $D$ and hence by $f_{\sigma}$.
@@
where $\sigma(1)$ is the set of edges in the $1$-skeleton of $\sigma$, and $\widetilde{u},\widetilde{w}\in\Sigma(0)$ are vertices connected by an edge in $\Sigma(1)$ that project to $u,w\in\sigma(0)$. Since $f$ is $\Gamma$-equivariant and the action of $\Gamma$ on $X$ is by isometries, $d(f(\widetilde{u}),f(\widetilde{w}))$ does not depend on the choice of $\widetilde{u},\widetilde{w}$.
@@
Since a $\Gamma$-equivariant function $f$ is determined by the associated function $f_{\sigma}:\sigma(0)\rightarrow X$, we can consider the energy functional as a continuous function $E:X^{n}\rightarrow\mathbb{R}$, where $n$ is the cardinality of $\sigma(0)$. The function $E$ is bounded below by zero, and Proposition 2.6 is equivalent to $E$ having a minimum. Therefore, if $I\geq 0$ is the infimum of $E$ in $X^{n}$, we need to show that $E(x)=I$ for some $x\in X^{n}$.
As the product of CAT(0) spaces is still CAT(0), so is the space $X^{n}$. Moreover, since the space $X$ has finite telescopic dimension and this property is preserved by taking products, the space $X^{n}$ has finite telescopic dimension. We refer to [caply] for more information and results on finite telescopic dimension for CAT(0) spaces.
For all $m\in\mathbb{N}$, let $S_{m}:=\{x\in X^{n}:\ E(x)\leq I+1/m\}$. Since the energy functional is convex (the distance function in a CAT(0) space is convex), the sets $(S_{m})_{m\in\mathbb{N}}$ are a descending chain of closed convex subsets of $X^{n}$. It is proved in [caply] (Theorem 1.1) that in this case either the intersection $\bigcap_{m\in\mathbb{N}}{S_{m}}$ is non-empty or the intersection of the visual boundaries $\bigcap_{m\in\mathbb{N}}{\partial S_{m}}$ is a non-empty subset of $\partial X$.
If the intersection $\bigcap_{m\in\mathbb{N}}{S_{m}}$ is non-empty, any $x$ in this intersection must satisfy $E(x)=I$ and hence we are done. In the remaining case, let $\xi\in\partial X$ be in the intersection $\bigcap_{m\in\mathbb{N}}{\partial S_{m}}$. To complete the proof of Proposition 2.6, we show:
@@
The point at infinity $\xi$ is fixed by the action of $\Gamma$.
@@
We can choose a sequence $(x_{m})_{m\in\mathbb{N}}$ of points in $X^{n}$ such that each $x_{m}$ is in $S_{m}$ and the sequence converges to $\xi$. For all $m\geq 1$, we have $E(x_{m})\leq I+1$.
Each point $x_{m}$ is associated with a $\Gamma$-equivariant function $f_{m}:\Sigma(0)\rightarrow X$. From the definition of the energy functional it follows that, whenever two vertices $v,w\in\Sigma(0)$ are connected by an edge in $\Sigma(1)$, we have $d(f_{m}(v),f_{m}(w))^{2}\leq E(f_{m})\leq I+1$.
Fix $g\in\Gamma$ and let $v_{1},\ldots,v_{n}$ be the vertices in the fundamental domain $D$. For every $m$, the point $x_{m}\in X^{n}$ has coordinates $(f_{m}(v_{1}),\ldots,f_{m}(v_{n}))$. From what we said above, for $1\leq i\leq n$ we have
$d(gf_{m}(v_{i}),f_{m}(v_{i}))=d(f_{m}(gv_{i}),f_{m}(v_{i}))\leq d_{\Sigma}(gv_{i},v_{i})\sqrt{I+1},$
where $d_{\Sigma}$ denotes the simplicial distance in $\Sigma$ (i. e. the graph distance in the $1$-skeleton). It follows that
$d(gx_{m},x_{m})\leq n\max_{1\leq i\leq n}{d_{\Sigma}(gv_{i},v_{i})}\sqrt{I+1}.$
This means that the distance $d(gx_{m},x_{m})$ is bounded above by a constant that does not depend on $m$. Therefore, since the sequence $(x_{m})_{m\in\mathbb{N}}$ converges to the point at infinity $\xi$, the sequence $(gx_{m})_{m\in\mathbb{N}}$ must also converge to $\xi$. This shows that $\xi$ is fixed by $g$.
@@ 2.3 Harmonic functions are constant
Therefore, we are left to prove Proposition 2.8. The proof is basically the same as in [wang]; we include it here in a form which is coherent with our situation and notation.
@@
We consider the variation at $v_{i}$ given by $y\in T_{f(v_{i})}X$. We have:
$\frac{d}{dt}E(t)$ $=\frac{d}{dt}\left(\sum_{u\in Lk(v_{i})(0)}{d(f(v_{i}),f(u))^{2}}\right)=$
$=\frac{d}{dt}\left(\sum_{u\in Lk(v_{i})(0)}{d(\mathrm{exp}(ty),f(u))^{2}}\right)=$
$=\sum_{u\in Lk(v_{i})(0)}{\langle y,-2Df|_{v_{i}}(u)\rangle},$
where the last equality holds because $f(v_{i})$ and $f(u)$ are connected by a minimal geodesic in $X$ and the gradient of the function $x\mapsto d(x,u)^{2}$ at $x=v_{i}$ is $-2Df|_{v_{i}}(u)$.
@@
We can now proceed with the proof of Proposition 2.8. Let $f$ be harmonic. Let $i\in\{1,\ldots,n\}$ and let $u,w\in Lk(v_{i})(0)$. Since $X$ is non-positively curved, we have
@@
Denoting the left-hand side by $||d(Df)_{v_{i}}||^{2}$ and summing over $i$, we obtain
$\sum_{i=1}^{n}{||d(Df)_{v_{i}}||^{2}}\leq\sum_{i=1}^{n}{\sum_{(u,w)\in Lk(v_{i})(1)}{d(f(u),f(w))^{2}}}.$ (1)
@@
$\sum_{i=1}^{n}{||d(Df)_{v_{i}}||^{2}}\leq(q+1)E(f).$ (2)
Consider now, for $1\leq i\leq n$, the function $g_{i}=Df|_{v_{i}}$. It is a function $g_{i}:Lk(v_{i})(0)\rightarrow T_{f(v_{i})}X$, and we know by Lemma 2.10 that $\sum_{u\in Lk(v_{i})(0)}{g_{i}(u)}=0$. We can apply Remark 2.3 and get $||dg_{i}||^{2}\geq\lambda||g_{i}||^{2}$ for some $\lambda>\frac{q+1}{2}$. This means that
@@
$\sum_{i=1}^{n}{||d(Df)_{v_{i}}||^{2}}\geq 2\lambda E(f).$ (3)
From (2) and (3) we get $(q+1)E(f)\geq 2\lambda E(f)$, which, since $\lambda>\frac{q+1}{2}$ and $E(f)\geq 0$, implies $E(f)=0$. By the definition of $E(f)$, this implies that $f$ is constant on $\Sigma(0)$, proving Proposition 2.8. As explained above, Theorem 2.1 follows.
@@ 3 Obtaining a fixed point inside $X$
In this section, we complete the proof of Theorem 1.4. We know from Theorem 2.1 that the action of the discrete torsion-free cocompact subgroup $\Gamma<G=\mathrm{SL}_{3}(\mathbb{F})$ on any $X\in\mathcal{X}$ has to fix a point in $\overline{X}$. For convenience, we name this property as in [wang]:
@@
If $\Gamma$ is $\sigma$-compact, Property (FH) is equivalent to Kazhdan’s Property (T). This is known as the Delorme-Guichardet Theorem ([kazhdanT], Theorem 2.12.4).
@@
Let $\Gamma$ be a locally compact group. Assume that every finite-index subgroup of $\Gamma$ has Property (F) and Property (FH). Then any continuous action by isometries of $\Gamma$ on a space $X\in\mathcal{X}$ has a fixed point in $X$.
@@ 3.1 More about the family $\mathcal{X}$
As a first step, we determine more explicitly the family of possible spaces $X$. On one hand, it contains all simply connected finite-dimensional symmetric spaces of non-positive curvature. On the other hand, it is shown in [duc2015] (Corollary 1.10) that any space in $\mathcal{X}$, after possibly removing the Euclidean de Rham factor, is isometric to a finite product of irreducible (finite-dimensional) symmetric spaces of non-compact type and spaces of the form $\mathrm{O}(p,\infty)/\mathrm{O}(p)\times\mathrm{O}(\infty)$, $\mathrm{U}(p,\infty)/\mathrm{U}(p)\times\mathrm{U}(\infty)$, $\mathrm{Sp}(p,\infty)/\mathrm{Sp}(p)\times\mathrm{Sp}(\infty)$ with $p\geq 1$, up to homothety. Such spaces were studied in [duc2013]. We use the notation $X_{p}(\mathbb{K})$ for these spaces, where $\mathbb{K}$ is the field of real, complex, or quaternionic numbers.
@@
We can reduce our problem to each of these single factors. According to a result by Foertsch and Lytchak [foly], every geodesic space $X$ of finite affine rank has a unique decomposition
@@
where $Y_{0}$ is a Euclidean space (possibly a point) and each $Y_{i}$, $1\leq i\leq n$, is an irreducible metric space not isometric to the real line, nor to a point. Moreover, each isometry of $X$ is a product of isometries of the single $Y_{i}$, up to possibly permuting factors which are isometric. This result was inspired by the classical De Rham decomposition for finite-dimensional Riemannian manifolds [derham].
@@
There is also an alternative way to see that the factors that appear in the decomposition (1) of $X$ coincide with the factors of Foertsch-Lytchak and therefore any isometry of $X$ is a product of isometries of the individual factors up to possibly permuting isometric ones. Indeed, since $X$ is a geodesic space and all its geodesics extend indefinitely in both directions, the same must be true for the factors $(Y_{i})_{0\leq i\leq n}$ in the Foertsch-Lytchak decomposition (this follows from the fact that geodesics in a product of metric spaces project to geodesics in factors, as seen in [bh], I.5.3). Then, as a consequence of Proposition 2.2 in [duc2023], these factors must be totally geodesic submanifolds of $X$ and therefore coincide (up to the order) with the factors in the decomposition (1). This was suggested by Bruno Duchesne.
Going back to the statement we wish to prove, Theorem 3.4, we see that $\Gamma$ must have a finite-index subgroup $\Gamma^{\prime}$ that acts on $X$ without permuting the factors. If we prove that the action of $\Gamma^{\prime}$ has a fixed point in $X$, then the action of $\Gamma$ will have a finite orbit and hence a fixed point ([bh], II.2.8). Therefore, from now on, we may just assume that the action of $\Gamma$ does not permute the factors.
Now, $\Gamma$ acts on each factor and the action on $X$ is just the product of these actions. The action on the Euclidean factor has a fixed point due to Property (FH). Each $M_{i}$ and each $N_{i}$ is a symmetric space that belongs to the class $\mathcal{X}$, hence the action of $\Gamma$ has a fixed point either on the space or on its boundary (Property (F)). If we prove Theorem 3.4 in the case where $X$ is a single factor, we will find that in the general case the action on each factor has a fixed point, and therefore the global action has a fixed point, proving the theorem. Hence from now on we will assume that $X$ is either an irreducible finite-dimensional symmetric space of non-compact type, or a space $X_{p}(\mathbb{K})$ for some $p\geq 1$ and $\mathbb{K}\in\{\mathbb{R},\mathbb{C},\mathbb{H}\}$.

### v2

@@ 2.2 Existence of harmonic functions
An approach with harmonic functions was also used in [18], where Duchesne proved a superrigidity result for cocompact lattices in higher rank semisimple Lie groups acting on spaces in the family $\mathcal{X}$.
Recall that $\Gamma$ is a discrete, torsion-free, cocompact subgroup of $G$. Its action on $\Sigma$ is free and proper ([24], Lemma 2.6). Therefore, $\Gamma$ is the fundamental group of the finite simplicial complex $\sigma=\Sigma/\Gamma$, and $\Gamma$-equivariant functions $f:\Sigma(0)\rightarrow X$ correspond to functions $f_{\sigma}:\sigma(0)\rightarrow X$, where $\sigma(0)$ is the set of vertices in $\sigma$. The correspondence is defined by choosing a fundamental domain $D$ of $\Gamma$ in $\Sigma(0)$ and assigning to $f_{\sigma}$ the values that $f$ takes at the vertices in $D$. Since $f$ is $\Gamma$-equivariant, all its values are uniquely determined by those at the vertices in $D$ and hence by $f_{\sigma}$.
@@
where $\sigma(1)$ is the set of edges in the $1$-skeleton of $\sigma$, and $\widetilde{u},\widetilde{w}\in\Sigma(0)$ are vertices connected by an edge in $\Sigma(1)$ that project to $u,w\in\sigma(0)$. Since $f$ is $\Gamma$-equivariant and the action of $\Gamma$ on $X$ is by isometries, $)d(f(\widetilde{u}),f(\widetilde{w}$)) does not depend on the choice of $\widetilde{u},\widetilde{w}$.
@@
Since a $\Gamma$-equivariant function $f$ is determined by its values on the fundamental domain $D$ for the $\Gamma$-action on $\Sigma(0)$, we can consider the energy functional as a continuous function $E:X^{n}\rightarrow\mathbb{R}_{\geq 0}$, where $n$ is the cardinality of $D$. The function $E$ is bounded below by zero, and Proposition 2.6 is equivalent to $E$ having a minimum. Therefore, if $I\geq 0$ is the infimum of $E$ in $X^{n}$, we need to show that $E(x)=I$ for some $x\in X^{n}$.
We are going to use a notion of dimension at large scale for our spaces $X\in\mathcal{X}$: the notion of telescopic dimension, which was defined in [12] as the supremum of geometric dimensions of asymptotic cones built from $X$. The notion of geometric dimension was defined inductively in [31] by setting the dimension of a discrete space to be 0 and defining $\dim(X)=\sup\{\dim(S_{x}X)+1|x\in X\}$, where $S_{x}X$ denotes the space of directions at the point $x$; this definition coincides with the supremum of topological dimensions of compact subsets in $X$.
It is proved in [17] (Corollary 1.10) that any space $X\in\mathcal{X}$ has finite telescopic dimension. As the product of CAT(0) spaces is still CAT(0), so is the space $X^{n}$; moreover, since the space $X$ has finite telescopic dimension and this property is preserved by taking products ([12], Lemma A.12), the space $X^{n}$ has finite telescopic dimension.
For all $m\in\mathbb{N}$, let $S_{m}:=\{x\in X^{n}:\ E(x)\leq I+1/m\}$. Since the energy functional is convex (the distance function in a CAT(0) space is convex), the sets $(S_{m})_{m\in\mathbb{N}}$ are a descending chain of closed convex subsets of $X^{n}$. It is proved in [12] (Theorem 1.1) that in this case either the intersection $\bigcap_{m\in\mathbb{N}}{S_{m}}$ is non-empty or the intersection of the visual boundaries $\bigcap_{m\in\mathbb{N}}{\partial S_{m}}$ is a non-empty subset of $\partial(X^{n})$ with intrinsic radius $\leq\pi/2$.
If the intersection $\bigcap_{m\in\mathbb{N}}{S_{m}}$ is non-empty, any $x$ in this intersection must satisfy $E(x)=I$ and hence we are done. Therefore, from now on, we can assume that $\bigcap_{m\in\mathbb{N}}{\partial S_{m}}$ is non-empty with intrinsic radius $\leq\pi/2$.
The action of $\Gamma$ on $X$ extends diagonally to an action on $X^{n}$. If we show that there exists $\xi\in\partial(X^{n})$ fixed by this action, then this implies that $\Gamma$ fixes a point in $\partial X$, since the boundary $(\partial X)^{n}$ is the spherical join of the boundaries of the factors ([9], II.9) and therefore $\xi$ corresponds to a $n$-tuple $(\xi_{1},\ldots,\xi_{n})\in(\partial X)^{n}$ together with weights $(\lambda_{1},\ldots,\lambda_{n})$ satisfying $\sum_{i=1}^{n}{\lambda_{i}^{2}}=1$, and $\xi_{i}\in\partial X$ must be fixed by $\Gamma$ whenever $\lambda_{i}\neq 0$. Therefore, to complete the proof of Proposition 2.6, it suffices to show:
@@
If the intersection $\bigcap_{m\in\mathbb{N}}{\partial S_{m}}$ is non-empty with intrinsic radius $\leq\pi/2$, then there exists $\xi\in\partial(X^{n})$ fixed by the action of $\Gamma$.
@@
Since $\Gamma$ acts by isometries on $X$, it is clear from Definition 2.4 that we have $E(gf)=E(f)$ for every $g\in\Gamma$ and for every $\Gamma$-equivariant function $f:\Sigma(0)\rightarrow X$. This means that every $\Gamma$ preserves the energy function $E$ on $X^{n}$ and therefore each of the energy level subsets $S_{m}\subset X^{n}$ is $\Gamma$-invariant. Therefore, each of the subsets $\partial S_{m}\subset\partial(X^{n})$ is also $\Gamma$-invariant, and the same is true for their intersection $\bigcap_{m\in\mathbb{N}}{\partial S_{m}}$. Since the CAT(0) space $X^{n}$ has finite telescopic dimension, its visual boundary $\partial(X^{n})$ endowed with the Tits metric has finite geometric dimension ([12], Proposition 2.1). Therefore, by ([3], Proposition 1.4), the $\Gamma$-invariant subset $\bigcap_{m\in\mathbb{N}}{\partial S_{m}}$ has a circumcenter that is fixed by $\Gamma$. This proof was inspired by a comment of a reviewer and by the proof of Proposition 1.8 in [12].
@@ 2.3 Harmonic functions are constant
Therefore, we are left to prove Proposition 2.8. In the case $G=\mathrm{SL}_{3}(\mathbb{F})$, the proof is basically the same as in [45]; we include it here in a form that is consistent with our situation and notation. In the case $G=\mathrm{Sp}_{4}(\mathbb{F})$, a small addition will be needed.
@@
We consider the variation at $v_{i}$ given by $y\in T_{f(v_{i})}X$. We denote by $E(t)$ the energy of the $\Gamma$-equivariant function that sends $v_{i}$ to $\mathrm{exp}(ty)$ and agrees with $f$ on $D\setminus\{v_{i}\}$. We have the following:
$\frac{d}{dt}E(t)=\frac{d}{dt}\left(\sum_{u\in Lk(v_{i})(0)}{d(\mathrm{exp}(ty),f(u))^{2}}\right)=\sum_{u\in Lk(v_{i})(0)}{\langle y,-2Df|_{v_{i}}(u)\rangle},$
where the last equality holds because $f(v_{i})$ and $f(u)$ are connected by a minimal geodesic in $X$ and the gradient of the function $x\mapsto d(x,u)^{2}$ at $x=v_{i}$ is $2Df|_{v_{i}}(u)$.
@@
We can now proceed with the proof of Proposition 2.8. First, we consider the case $G=\mathrm{SL}_{3}(\mathbb{F})$, where Proposition 2.2 holds for $\lambda(v)=q+1-\sqrt{q}$ that does not depend on $v$.
Let $f$ be harmonic. Let $i\in\{1,\ldots,n\}$ and let $u,w\in Lk(v_{i})(0)$. Since $X$ is non-positively curved, we have
@@
Denoting the left-hand side by $d(Df)_{v_{i}}||^{2}$ and summing over $i$, we obtain
[…]
@@
[…]
@@
[…]
From (2) and (3) we get $(q+1)E(f)\geq 2\lambda E(f)$. We have $\lambda=q+1-\sqrt{q}>\frac{q+1}{2}$ for $q\geq 2$ and since $E(f)\geq 0$, this implies $E(f)=0$. By the definition of $E(f)$, this means that $f$ is constant on $\Sigma(0)$, proving Proposition 2.8 in the case $G=\mathrm{SL}_{3}(\mathbb{F})$.
We are left with the case $G=\mathrm{Sp}_{4}(\mathbb{F})$. Here $\lambda(v)$ is not the same for every $v$, but depends on whether the vertex $v$ is special or not. However, we can repeat the same proof and sum separately over special and non-special vertices. Let $\{v_{1},\ldots,v_{n}\}=S\sqcup NS$ be the partition of $D=\{v_{1},\ldots,v_{n}\}$ into special and non-special vertices, respectively. With Definition 2.4 in mind, we can set $E(f)=E_{1}(f)+E_{2}(f)$, where $E_{1}(f)$ is defined by taking the sum only over edges between two special vertices, and $E_{2}(f)$ is defined by taking the sum only over edges between a special and a non-special vertex (recall that there is no edge between two non-special vertices).
We can obtain inequality (1) separately for the sum over $S$ and for the sum over $NS$. Since every chamber in the building has one non-special and two special vertices, inequality (2) becomes:
[…]
We can continue the proof as above, again summing separately over $S$ and over $NS$. Since $\lambda(v)=q+1-\sqrt{2q}$ for all $v\in S$ and $\lambda(v)=q+1$ for all $v\in NS$, inequality (3) becomes:
[…]
Putting this together with the previous inequalities, we get $E_{1}(f)\geq E_{2}(f)$ from the sum over $NS$, and then $(q+1)E_{2}(f)\geq 3(q+1-\sqrt{2q})E_{2}(f)$ from the sum over $S$. If $E_{2}(f)>0$, this is only possible if $\frac{\sqrt{2q}}{q+1}\geq\frac{2}{3}$, which implies $q\leq 2$. This means that if $q\geq 3$, we must have $E_{2}(f)=0$ and therefore $f$ constant, completing the proof of Proposition 2.8. As explained above, Theorem 2.1 follows.
Remark 2.11.
The last part of the proof of Proposition 2.8 is the reason why in Theorem 1.1 we need to assume that the residue field of $\mathbb{F}$ has at least three elements. In Theorem 1.2 we do not need this assumption, since in the case $G=\mathrm{SL}_{3}(\mathbb{F})$ the inequality works for all $q\geq 2$.
@@ 3 Obtaining a fixed point inside $X$
In this section, we complete the proof of Theorem 1.6. We know from Theorem 2.1 that the action of the torsion-free cocompact lattice $\Gamma<G$ on any $X\in\mathcal{X}$ has to fix a point in $\overline{X}$. For convenience, we name this property as in [45]:
@@
If $\Gamma$ is $\sigma$-compact, Property (FH) is equivalent to Kazhdan’s Property (T). This is known as the Delorme-Guichardet Theorem ([4], Theorem 2.12.4).
@@
Let $\Gamma$ be a locally compact group. Assume that $\Gamma$ has Property (F) and Property (FH). Then any continuous action by isometries of $\Gamma$ on a space $X\in\mathcal{X}$ has a fixed point in $X$.
Remark 3.5.
If $\Gamma$ has Property (F) or Property (FH), every closed finite index subgroup $\Gamma^{\prime}<\Gamma$ has the corresponding property.
Proof of Remark 3.5.
For Property (FH), this is ([4], Proposition 2.5.7). For Property (F), the proof is similar. If $\Gamma^{\prime}$ has a continuous action $\alpha$ by isometries on a space $X\in\mathcal{X}$, one can define the induced action $\mathrm{Ind}_{\Gamma^{\prime}}^{\Gamma}{\alpha}$ on the space $\Xi$ of functions $\xi:G\rightarrow X$ such that $\xi(gh)=\alpha(h^{-1})\xi(g)$, by
$\left(\mathrm{Ind}_{\Gamma^{\prime}}^{\Gamma}{\alpha}(g)\xi\right)(x)=\xi(g^{-1}x),\ g,x\in\Gamma,\ \xi\in\Xi.$
This construction is a special case of the one that appears in ([34], 5.4).
Let $n=[\Gamma:\Gamma^{\prime}]$. We can identify the space $\Xi$ with $X^{n}$ by choosing coset representatives, and it is easy to check that the action $\mathrm{Ind}_{\Gamma^{\prime}}^{\Gamma}{\alpha}$ is continuous and isometric for the $L^{2}$ product metric. We have $\Xi\in\mathcal{X}$, therefore the action fixes a point in $\xi\in\Xi$ or $\eta\in\partial\Xi$. In the first case, it is immediately checked that $\xi(e)\in X$ is fixed by $H$. In the second case, we know that $\partial\Xi$ is the spherical join of $n$ copies of $\partial X$ ([9], II.9), that is, the space of $n$-tuples $(\zeta_{1},\ldots,\zeta_{n})\in(\partial X)^{n}$ together with weights $(\lambda_{1},\ldots,\lambda_{n})$ satisfying $\sum_{i=1}^{n}{\lambda_{i}^{2}}=1$. It is immediately checked that $\eta$ must correspond to a $n$-tuple of the form $(\zeta,\ldots,\zeta)$ with equal weights, with $\zeta\in\partial X$ fixed by $H$.
@@ 3.1 More about the family $\mathcal{X}$
As a first step towards the proof of Theorem 3.4, we determine more explicitly the family of possible spaces $X$. On one hand, it contains all simply connected finite-dimensional symmetric spaces of non-positive curvature. On the other hand, it is shown in [17] (Corollary 1.10) that any space in $\mathcal{X}$, after possibly removing the Euclidean de Rham factor, is isometric to a finite product of irreducible (finite-dimensional) symmetric spaces of non-compact type and spaces of the form $\mathrm{O}(p,\infty)/\mathrm{O}(p)\times\mathrm{O}(\infty)$, $\mathrm{U}(p,\infty)/\mathrm{U}(p)\times\mathrm{U}(\infty)$, $\mathrm{Sp}(p,\infty)/\mathrm{Sp}(p)\times\mathrm{Sp}(\infty)$ with $p\geq 1$, up to homothety. Such spaces were studied in [16]. We use the notation $X_{p}(\mathbb{K})$ for these spaces, where $\mathbb{K}$ is the field of real, complex, or quaternionic numbers.
@@
We can reduce our problem to each of these single factors. According to a result by Foertsch and Lytchak [23], every geodesic space $X$ of finite affine rank has a unique decomposition
@@
where $Y_{0}$ is a Euclidean space (possibly a point) and each $Y_{i}$, $1\leq i\leq n$, is an irreducible metric space not isometric to the real line or to a point. Moreover, each isometry of $X$ is a product of isometries of the single $Y_{i}$, up to possibly permuting factors which are isometric. This result is a generalization of the classical De Rham decomposition for finite-dimensional Riemannian manifolds [40].
@@
There is also an alternative way to see that the factors that appear in the decomposition (1) of $X$ coincide with the factors of Foertsch-Lytchak and therefore any isometry of $X$ is a product of isometries of the individual factors up to possibly permuting isometric ones. Indeed, since $X$ is a geodesic space and all its geodesics extend indefinitely in both directions, the same must be true for the factors $(Y_{i})_{0\leq i\leq n}$ in the Foertsch-Lytchak decomposition (this follows from the fact that geodesics in a product of metric spaces project to geodesics in factors, as seen in [9], I.5.3). Then, as a consequence of Proposition 2.2 in [19], these factors must be totally geodesic submanifolds of $X$ and therefore coincide (up to the order) with the factors in the decomposition (1). This was suggested by Bruno Duchesne.
Going back to the statement we wish to prove, Theorem 3.4, we see that $\Gamma$ must have a finite index subgroup $\Gamma^{\prime}$ that acts on $X$ without permuting the factors. If we prove that the action of $\Gamma^{\prime}$ has a fixed point in $X$, then the action of $\Gamma$ will have a finite orbit and hence a fixed point ([9], II.2.8). Therefore, from now on, we may just assume that the action of $\Gamma$ does not permute the factors.
Now, $\Gamma$ acts on each factor and the action on $X$ is just the product of these actions. The action on the Euclidean factor has a fixed point due to Property (FH). Each $M_{i}$ and each $N_{i}$ is a symmetric space that belongs to the class $\mathcal{X}$, hence the action of $\Gamma$ has a fixed point in the space or in its boundary (Property (F)). If we prove Theorem 3.4 in the case where $X$ is a single factor, we will find that in the general case the action on each factor has a fixed point, and therefore the global action has a fixed point, proving the theorem. Hence from now on we will assume that $X$ is either an irreducible finite-dimensional symmetric space of non-compact type, or a space $X_{p}(\mathbb{K})$ for some $p\geq 1$ and $\mathbb{K}\in\{\mathbb{R},\mathbb{C},\mathbb{H}\}$.